# HiCL: Hippocampal-Inspired Continual Learning

### v1

@@ Abstract
We propose HiCL, a novel hippocampal-inspired dual-memory continual learning architecture designed to mitigate catastrophic forgetting by using elements inspired by the hippocampal circuitry. Our system encodes inputs through a grid-cell-like layer, followed by sparse pattern separation using a dentate gyrus-inspired module with top-$k$ sparsity. Episodic memory traces are maintained in a CA3-like autoassociative memory. Task-specific processing is dynamically managed via a DG-gated mixture-of-experts mechanism, wherein inputs are routed to experts based on cosine similarity between their normalized sparse DG representations and learned task-specific DG prototypes computed through online exponential moving averages. This biologically grounded yet mathematically principled gating strategy enables differentiable, scalable task-routing without relying on a separate gating network, and enhances the model’s adaptability and efficiency in learning multiple sequential tasks. Cortical outputs are consolidated using Elastic Weight Consolidation weighted by inter-task similarity. Crucially, we incorporate prioritized replay of stored patterns to reinforce essential past experiences. Evaluations on standard continual learning benchmarks demonstrate the effectiveness of our architecture in reducing task interference, achieving near state-of-the-art results in continual learning tasks at lower computational costs.
@@ Introduction
In continual learning, every new task is like a distinct episode in our life’s unfolding story. Just as the hippocampus rapidly encodes unique experiences into episodic memory, tagging them with contextual and temporal cues, a continual learner must capture each task’s defining features before they fade. The brain then replays these episodes—during sleep or quiet wakefulness—to gradually integrate them into neocortical networks, preventing older memories from being overwritten by new ones. We leverage this complementary learning systems principle by treating each task as an episodic trace: we store compact representations in a replay buffer and periodically revisit them during training, mirroring how hippocampal replay consolidates previous and precious episodes into lasting knowledge [1].
Figure 1: Schematic of the HiCl architecture overlaid with the hippocampal trisynaptic circuit. The entorhinal cortex (layers II/III) projects to the dentate gyrus (DG), which performs pattern separation. DG outputs feed into CA3, an autoassociative network responsible for pattern completion. CA3 then projects to CA1, which integrates inputs from both CA3 and the entorhinal cortex before sending output to other brain regions. This circuit underlies memory encoding and retrieval processes. The general background here of the hippocampal circuit is based on [2] and adapted from [3, 4, 5, 6].
@@
Dentate Gyrus (DG): In biology, the DG performs pattern separation by mapping inputs onto a very sparse, typically only $3–5\%$ of granule cells fire for any given pattern [7]. In our model, the SparseActivation DG layer enforces top‑$k$ sparsity $(k=5\%)$, orthogonalizing features both for expert representations and for routing signals.
CA3: The DG’s orthogonal codes feed into CA3’s recurrent attractor network, which functions as pattern completion, reconstructing full memory traces from partial cues [8]. We mirror this with a lightweight two-layer MLP that refines and transforms DG outputs via a non-linear projection, functionally completing the pattern for downstream processing.
CA1: Finally, CA1 integrates completed patterns from CA3 with direct cortical inputs, acting as a comparator and coordinating memory consolidation through hippocampo‑cortical dialogue [9]. We emulate this via our ca1_integration block and a consolidation stage combining Elastic Weight Consolidation (EWC) with a replay buffer.
@@
This biology‑inspired solution to the stability–plasticity dilemma offers actionable insights for continual learning in artificial systems. Our DG‑Gated Mixture‑of‑Experts architecture translates hippocampal principles into three core mechanisms: (1) a top-$k$ sparse DG layer ($k\approx 5\%$) that orthogonalizes inputs for both expert encoding and dynamic routing; (2) a CA3‑inspired two‑layer MLP autoencoder that completes partial patterns via attractor-like reconstruction; and (3) a consolidation module combining a prioritized replay buffer with Elastic Weight Consolidation to protect prior knowledge.
By instantiating the hippocampal tripartite circuit across multiple experts and leveraging DG‑based gating to minimize interference, CA3‑style pattern completion for robust recall, and replay+EWC for selective consolidation, our DG‑Gated MoE exemplifies how neuroscience‑informed architectures can achieve efficient, adaptive continual learning.
@@
The CLS theory posits a dual-memory system: a fast hippocampal module for episodic learning and a slow neocortical system for generalization [1, 10]. This foundational concept has influenced many biologically inspired CL frameworks. Examples include hybrid cortico-hippocampal networks that combine feedforward and spiking layers [11], DualNet’s distinct fast/slow learners [12], FearNet’s use of generative replay [13], and Triple Memory Networks that model hippocampus-PFC interactions with replay and consolidation [14].
@@
Replay of past experiences is a powerful anti-forgetting mechanism, mirroring the brain’s hippocampal replay during sleep and wakefulness [15]. In machine learning, this is realized through experience replay, which explicitly rehearses stored examples [16], or generative replay, which trains a network to synthesize pseudo-data [17]. We adopt prioritized experience replay, inspired by findings that emphasizing important memories enhances learning efficiency in both reinforcement learning [18] and biological systems [19]. The ability to generate contrastive data during replay is considered crucial for preventing representational collapse [20].
@@
Many CL frameworks are categorized as parameter isolation, rehearsal-based, or regularization-based [21].
  • Regularization methods add a loss penalty to prevent changes to weights important for past tasks, as seen in Elastic Weight Consolidation (EWC) [22], Learning without Forgetting (LwF) [23], and Synaptic Intelligence (SI) [24].
  • Parameter isolation methods assign distinct weights to different tasks, using techniques like network expansion in Progressive Nets [25] or hard attention masks [26, 27, 28].
  • Rehearsal methods use a buffer of past data [29, 30, 31] or a generative model [17, 32, 33]. Hybrid models like Progress & Compress [34] and CLEAR [30] combine these strategies effectively. Our approach synergizes these ideas, using architectural modularity, sparsity, and replay.
@@
MoE models enable modular specialization by routing inputs to specialized subnetworks [35, 36], a principle recently applied to CL [37, 38] and large-scale models [39, 40]. Stability is often improved via adaptive routing, for instance by annealing a routing temperature [41, 42]. Our DG-Gated MoE is distinct in that expert selection emerges via cosine similarity to prototypes, bypassing an explicit learned gating network.
@@
Storing class prototypes instead of raw data is a memory-efficient rehearsal strategy pioneered by iCaRL [29]. Subsequent work has combined prototype anchoring with contrastive objectives to preserve the geometry of the feature space across tasks [43, 44]. We follow this line, maintaining online exponential moving average (EMA) prototypes and applying a push-pull contrastive loss during replay.
@@
Our design is grounded in the hippocampal trisynaptic circuit (DG $\rightarrow$ CA3 $\rightarrow$ CA1) and the theory of complementary learning systems (CLS) [1, 10]. We operationalize these motifs as inductive scaffolds: structured encoding, sparse separation, nonlinear refinement/completion, and integration, which together feed into prototype-based routing and consolidation. Replay-inspired mechanisms and prototype anchoring provide retention and interference mitigation, while modular gating arises naturally from the quasi-orthogonal codes produced by DG-like separation. Learning is organized as a two-phase dual-memory schedule—specialization followed by consolidation—whose detailed losses are defined in the Architecture section.
@@
Grid-cell-like representations from the entorhinal cortex provide a structured relational prior that has been shown to enhance spatial and contextual reasoning in downstream tasks [45, 46]. We implement this by applying $M=4$ parallel $1\times 1$ convolutions with learned phase offsets to backbone features $\mathbf{f}$:
@@
To reduce interference between tasks, we draw on the role of the dentate gyrus in orthogonalizing inputs via sparse coding [4, 47]. The encoded representation $\mathbf{g}$ is projected and sparsified:
@@
This top-$k$ mechanism approximates inhibitory microcircuitry to produce quasi-orthogonal sparse codes, creating natural modularization and enabling routing without explicit task labels, akin to capacity partitioning in MoE and gating literature [35, 36].
@@
The sparse separation codes are refined through a feed-forward MLP to approximate the pattern completion function attributed to CA3 [48]. Unlike classical recurrent attractor implementations [49], we use:
@@
providing nonlinear task-specific refinement while maintaining stability in the face of non-i.i.d. data [50, 51].
@@
This echoes prototype-based rehearsal approaches [29, 43] and supplies stable references for both gating and later consolidation. The use of replay-like alignment and contrastive objectives (detailed in Section The Architecture) is inspired by hippocampal replay phenomena that reinforce important memories during offline phases [15, 18, 19].
From Motifs to Mechanism.
The preceding motifs define the representations and anchors—grid-cell priors, sparse pattern-separated codes, refined completions, and maintained prototypes—that the model builds on. The next section gives the concrete architectural instantiation, including how these pieces are composed into a DG-Gated Mixture-of-Experts and how the two-phase dual-memory objectives are applied.
@@
$\mathcal{L}_{\mathrm{intra}}$ $=\sum_{i,j}\mathbf{1}_{[y_{i}=y_{j}]}\|\mathbf{p}_{i}-\mathbf{p}_{j}\|^{2}$ (2)
@@
Thus:
$\mathcal{L}_{\mathrm{phase1}}$ $=\mathcal{L}_{\mathrm{cls}}+\alpha_{\mathrm{intra}}\,\mathcal{L}_{\mathrm{intra}}+\alpha_{\mathrm{rep}}\,\mathcal{L}_{\mathrm{replay}}$
$\quad+\alpha_{\mathrm{dist}}\,\mathcal{L}_{\mathrm{distill}}+\alpha_{\mathrm{ewc}}\,\mathcal{L}_{\mathrm{EWC}}+\alpha_{s}\,\mathcal{L}_{\mathrm{sparsity}}.$ (4)
@@ Results
TABLE I: Comparison of accuracy (%), and FLOPs (M) on Split CIFAR-10 and Split TinyImageNet under different memory and compute settings. The baselines have been updated based on new results for Split CIFAR-10.
Method Buffer Sparsity Split CIFAR-10 Split TinyImageNet
Size Task-IL ↑ Class-IL ↑ FLOPs (M) ↓ Task-IL ↑ Class-IL ↑ FLOPs (M) ↓
@@
HiCL Small 200 0.15 $90.24\pm 0.84$ $57.65\pm 0.49$ 38.21 $59.16\pm 0.37$ $23.28\pm 0.46$ 152.9
500 0.15 $90.36\pm 0.54$ $60.68\pm 1.05$ 38.21 $59.63\pm 0.85$ $23.51\pm 0.91$ 152.9
HiCL Large 200 0.15 $91.74\pm 0.81$ $60.48\pm 0.15$ 84.18 $\textbf{62.76}\pm 0.26$ $25.45\pm 0.14$ 336.71
500 0.15 $92.38\pm 0.49$ $65.58\pm 0.45$ 84.18 $62.60\pm 0.79$ $\textbf{25.71}\pm 0.64$ 336.71
@@
  • Split CIFAR-10 [52]: The 10 classes are divided into 5 sequential tasks, with 2 new classes introduced in each task.
  • Split Tiny-ImageNet [53]: A more complex benchmark where 200 classes are split into 10 tasks of 20 classes each.
@@
We use a 2‐conv + 2‐FC LeNet backbone for CIFAR‐10 and a deeper CNN for Tiny‐ImageNet. The GridCellLayer has $M=4$ sinusoidal units, DG size =1024=1024 with top‐$k$ sparsity $k=5\%$, CA3/CA1 MLP widths $(512,256)$ and $(512,256,128)$ respectively, and replay buffer of 200 samples per task. Models are trained with Adam for 10 epochs per CIFAR‐10 task (20 for Tiny‐ImageNet), $\lambda_{\mathrm{ewc}}=0.1$, and contrastive margin $m=0.2$.
Table I reports Task‑IL accuracy, Class‑IL accuracy, and inference FLOPs per forward pass for both Small and Large HiCL variants alongside a suite of baselines. We fix the buffer at 200 or 500 samples per task and evaluate on Split CIFAR‑10 (columns 4–6) and Split Tiny‑ImageNet (columns 7–9). Baselines are grouped by methodology: regularization‑based, replay‑based, architectural/hybrid, and the joint offline upper bound.

### v2

@@ Abstract
We propose HiCL, a novel hippocampal-inspired dual-memory continual learning architecture designed to mitigate catastrophic forgetting by using elements inspired by the hippocampal circuitry. Our system encodes inputs through a grid-cell-like layer, followed by sparse pattern separation using a dentate gyrus-inspired module with top-$k$ sparsity. Episodic memory traces are maintained in a CA3-like autoassociative memory. Task-specific processing is dynamically managed via a DG-gated mixture-of-experts mechanism, wherein inputs are routed to experts based on cosine similarity between their normalized sparse DG representations and learned task-specific DG prototypes computed through online exponential moving averages. This biologically grounded yet mathematically principled gating strategy enables differentiable, scalable task-routing without relying on a separate gating network, and enhances the model’s adaptability and efficiency in learning multiple sequential tasks. Cortical outputs are consolidated using Elastic Weight Consolidation weighted by inter-task similarity. Crucially, we incorporate prioritized replay of stored patterns to reinforce essential past experiences. Evaluations on standard continual learning benchmarks demonstrate the effectiveness of our architecture in reducing task interference, achieving near state-of-the-art results in continual learning tasks at lower computational costs. Our code is available here https://github.com/kushalk173-sc/HiCL.
@@ Introduction
In continual learning, every new task is like a distinct episode in our life’s unfolding story. Just as the hippocampus rapidly encodes unique experiences into episodic memory, tagging them with contextual and temporal cues, a continual learner must capture each task’s defining features before they fade. The brain then replays these episodes—during sleep or quiet wakefulness—to gradually integrate them into neocortical networks, preventing older memories from being overwritten by new ones. We leverage this complementary learning systems principle by treating each task as an episodic trace: we store compact representations in a replay buffer and periodically revisit them during training, mirroring how hippocampal replay consolidates previous and precious episodes into lasting knowledge (McClelland, McNaughton, and O’Reilly 1995).
Figure 1: Schematic of the HiCl architecture overlaid with the hippocampal trisynaptic circuit. The entorhinal cortex (layers II/III) projects to the dentate gyrus (DG), which performs pattern separation. DG outputs feed into CA3, an autoassociative network responsible for pattern completion. CA3 then projects to CA1, which integrates inputs from both CA3 and the entorhinal cortex before sending output to other brain regions. This circuit underlies memory encoding and retrieval processes. The general background here of the hippocampal circuit is based on (Petrantonakis and Poirazi 2014) and adapted from Moser, Kropff, and Moser (2008); Leutgeb et al. (2007); Rolls (2013b); Otmakhova, Otmakhov, and Griffith (2018).
@@
Dentate Gyrus (DG): In biology, the DG performs pattern separation by mapping inputs onto a very sparse, typically only $3–5\%$ of granule cells fire for any given pattern (Hainmueller and Bartos 2020). In our model, the SparseActivation DG layer enforces top‑$k$ sparsity $(k=5\%)$, orthogonalizing features both for expert representations and for routing signals.
CA3: The DG’s orthogonal codes feed into CA3’s recurrent attractor network, which functions as pattern completion, reconstructing full memory traces from partial cues (Gilbert and Brushfield 2009). We mirror this with a lightweight two-layer MLP that refines and transforms DG outputs via a non-linear projection, functionally completing the pattern for downstream processing.
CA1: Finally, CA1 integrates completed patterns from CA3 with direct cortical inputs, acting as a comparator and coordinating memory consolidation through hippocampo‑cortical dialogue (Bartsch et al. 2011). We emulate this via our ca1_integration block and a consolidation stage combining Elastic Weight Consolidation (EWC) with a replay buffer.
@@
This biology‑inspired solution to the stability–plasticity dilemma offers actionable insights for continual learning in artificial systems. Our DG‑Gated Mixture‑of‑Experts architecture translates hippocampal principles into three core mechanisms: (1) a top-$k$ sparse DG layer ($k\approx 5\%$) that orthogonalizes inputs for both expert encoding and dynamic routing; (2) a CA3‑inspired two‑layer MLP autoencoder that completes partial patterns via attractor-like reconstruction; and (3) a consolidation module combining a prioritized replay buffer with Elastic Weight Consolidation to protect prior knowledge. By instantiating the hippocampal tripartite circuit across multiple experts and leveraging DG‑based gating to minimize interference, CA3‑style pattern completion for robust recall, and replay+EWC for selective consolidation, our DG‑Gated MoE exemplifies how neuroscience‑informed architectures can achieve efficient, adaptive continual learning.
@@
The CLS theory posits a dual-memory system: a fast hippocampal module for episodic learning and a slow neocortical system for generalization (McClelland, McNaughton, and O’Reilly 1995; Kemker et al. 2018). This foundational concept has influenced many biologically inspired CL frameworks. Examples include hybrid cortico-hippocampal networks that combine feedforward and spiking layers (Rothschild, Eban, and Frank 2017), DualNet’s distinct fast/slow learners (Pham, Liu, and Hoi 2021), FearNet’s use of generative replay (Kemker and Kanan 2017), and Triple Memory Networks that model hippocampus-PFC interactions with replay and consolidation (Wang et al. 2021).
@@
Replay of past experiences is a powerful anti-forgetting mechanism, mirroring the brain’s hippocampal replay during sleep and wakefulness (Cairney et al. 2018). In machine learning, this is realized through experience replay, which explicitly rehearses stored examples (Rolnick et al. 2019), or generative replay, which trains a network to synthesize pseudo-data (Shin et al. 2017). We adopt prioritized experience replay, inspired by findings that emphasizing important memories enhances learning efficiency in both reinforcement learning (Schaul et al. 2016) and biological systems (Mattar and Daw 2018). The ability to generate contrastive data during replay is considered crucial for preventing representational collapse (Hinton 2022).
@@
Many CL frameworks are categorized as parameter isolation, rehearsal based, or regularization based (De Lange et al. 2021). Regularization methods add a loss penalty to prevent changes to weights important for past tasks, as seen in Elastic Weight Consolidation (EWC) (Kirkpatrick et al. 2017), Learning without Forgetting (LwF) (Li and Hoiem 2017), and Synaptic Intelligence (SI) (Zenke, Poole, and Ganguli 2017). Parameter isolation methods assign distinct weights to different tasks, using techniques such as network expansion in Progressive Nets (Rusu et al. 2016) or hard attention masks (Serrà et al. 2018; Madotto et al. 2021; Wang et al. 2024). Rehearsal methods use a buffer of past data (Rebuffi et al. 2017; Riemer et al. 2018; Zhang, Wang, and Yang 2022) or a generative model (Shin et al. 2017; Su et al. 2019; Skierś and Deja 2025). Hybrid models such as Progress & Compress (Schwarz et al. 2018) and CLEAR (Riemer et al. 2018) combine these strategies effectively, and our approach synergizes these ideas by using architectural modularity, sparsity, and replay.
@@
MoE models enable modular specialization by routing inputs to specialized subnetworks (Yuksel, Wilson, and Gader 2012; Shazeer et al. 2017), a principle recently applied to CL (Li et al. 2025; Le et al. 2024) and large-scale models (Guo et al. 2025; Cai et al. 2025). Stability is often improved via adaptive routing, for instance by annealing a routing temperature (Nie et al. 2022; Fedus, Zoph, and Shazeer 2022). Our DG-Gated MoE is distinct in that expert selection emerges via cosine similarity to prototypes, bypassing an explicit learned gating network.
@@
Storing class prototypes instead of raw data is a memory-efficient rehearsal strategy pioneered by iCaRL (Rebuffi et al. 2017). Subsequent work has combined prototype anchoring with contrastive objectives to preserve the geometry of the feature space across tasks (Caccia et al. 2021; Aghasanli, Li, and Angelov 2025). We follow this line, maintaining online exponential moving average (EMA) prototypes and applying a push-pull contrastive loss during replay.
@@
Our design is grounded in the hippocampal trisynaptic circuit (DG $\rightarrow$ CA3 $\rightarrow$ CA1) and the theory of complementary learning systems (CLS) (McClelland, McNaughton, and O’Reilly 1995; Kemker et al. 2018). We operationalize these motifs as inductive scaffolds: structured encoding, sparse separation, nonlinear refinement/completion, and integration, which together feed into prototype-based routing and consolidation. Replay-inspired mechanisms and prototype anchoring provide retention and interference mitigation, while modular gating arises naturally from the quasi-orthogonal codes produced by DG-like separation. Learning is organized as a two-phase dual-memory schedule—specialization followed by consolidation—whose detailed losses are defined in the Architecture section.
@@
Grid-cell-like representations from the entorhinal cortex provide a structured relational prior that has been shown to enhance spatial and contextual reasoning in downstream tasks (Hafting et al. 2005; Banino et al. 2018). We implement this by applying $M=4$ parallel $1\times 1$ convolutions with learned phase offsets to backbone features $\mathbf{f}$:
@@
To reduce interference between tasks, we draw on the role of the dentate gyrus in orthogonalizing inputs via sparse coding (Leutgeb et al. 2007; Rolls 2013a). The encoded representation $\mathbf{g}$ is projected and sparsified:
@@
This top-$k$ mechanism approximates inhibitory microcircuitry to produce quasi-orthogonal sparse codes, creating natural modularization and enabling routing without explicit task labels, akin to capacity partitioning in MoE and gating literature (Yuksel, Wilson, and Gader 2012; Shazeer et al. 2017).
@@
The sparse separation codes are refined through a feed-forward MLP to approximate the pattern completion function attributed to CA3 (Treves and Rolls 1994). Unlike classical recurrent attractor implementations (Hopfield 1982), we use:
@@
providing nonlinear task-specific refinement while maintaining stability in the face of non-i.i.d. data (Alonso and Krichmar 2024; Krotov and Hopfield 2016).
@@
Biological Abstraction.
Our design philosophy emphasizes principled abstraction, extracting core computational principles from hippocampal circuits while trading biological fidelity for computational efficiency and trainability. The biological DG employs complex inhibitory microcircuitry to enforce sparse coding (Hainmueller and Bartos 2020). We replace this with a non-parametric TopK operation that, when combined with Phase II contrastive tuning, achieves effective pattern separation without simulating neuronal dynamics. Similarly, CA3’s recurrent attractor network (Treves and Rolls 1994) is approximated by a feedforward MLP that captures pattern completion functionality in a single stable forward pass, avoiding the computational cost and training instability of recurrent architectures (Hopfield 1982). Our learnable GridCellLayer adapts its spatial encoding to dataset geometry rather than using fixed Fourier bases, providing representational flexibility for the continual learning setting. These design choices demonstrate that capturing biological function enables effective continual learning without requiring literal anatomical replication.
@@
This echoes prototype-based rehearsal approaches (Rebuffi et al. 2017; Caccia et al. 2021) and supplies stable references for both gating and later consolidation. The use of replay-like alignment and contrastive objectives (detailed in Section The Architecture) is inspired by hippocampal replay phenomena that reinforce important memories during offline phases (Cairney et al. 2018; Schaul et al. 2016; Mattar and Daw 2018).
@@
Each expert maintains its own DG module, enabling task-specific sparse representations. At inference, inputs are processed through all $N$ DG modules in parallel to compute routing scores via cosine similarity to prototypes. Once the expert is selected, only that expert’s CA3 and CA1 modules are activated for prediction. The gating computation requires $N$ DG forward passes, included in our reported FLOPs.
@@
$\mathcal{L}_{\mathrm{intra}}$ $=\sum_{i,j}\mathbf{1}_{[y_{i}=y_{j}]}\|\mathbf{p}_{i}-\mathbf{p}_{j}\|^{2}$
@@
The Phase I objective combines these terms:
$\mathcal{L}_{\mathrm{phase1}}$ $=\mathcal{L}_{\mathrm{cls}}+\alpha_{\mathrm{intra}}\,\mathcal{L}_{\mathrm{intra}}+\alpha_{\mathrm{dist}}\,\mathcal{L}_{\mathrm{distill}}+\alpha_{s}\,\mathcal{L}_{\mathrm{sparsity}}.$ (3)
@@ Results
Method Buffer Split CIFAR 10 Split TinyImageNet
Size Task IL ↑ Class IL ↑ FLOPs (M) ↓ Task IL ↑ Class IL ↑ FLOPs (M) ↓
@@
HiCL Small 200 $90.24\pm 0.84$ $57.65\pm 0.49$ 38.21 $59.16\pm 0.37$ $23.28\pm 0.46$ 152.9
(Sparsity: 0.15) 500 $90.36\pm 0.54$ $60.68\pm 1.05$ 38.21 $59.63\pm 0.85$ $23.51\pm 0.91$ 152.9
HiCL Large 200 $91.74\pm 0.81$ $60.48\pm 0.15$ 84.18 $\textbf{62.76}\pm 0.26$ $25.45\pm 0.14$ 336.71
(Sparsity: 0.15) 500 $92.38\pm 0.49$ $65.58\pm 0.45$ 84.18 $62.60\pm 0.79$ $\textbf{25.71}\pm 0.64$ 336.71
Table 1: Comparison of accuracy (%), and FLOPs (M) on Split CIFAR-10 and Split TinyImageNet under different memory and compute settings. The baselines have been updated based on new results for Split CIFAR-10.
@@
  • Split CIFAR-10 (Krizhevsky, Hinton et al. 2009): The 10 classes are divided into 5 sequential tasks, with 2 new classes introduced in each task.
  • Split Tiny-ImageNet (Deng et al. 2009): A more complex benchmark where 200 classes are split into 10 tasks of 20 classes each.
@@
We use a 2‐conv + 2‐FC LeNet backbone for CIFAR‐10 and and a 4-conv + 2-FC CNN for Tiny-ImageNet (conv layers: 64→128→256→512 filters with 3×3 kernels and max pooling). The GridCellLayer has $M=4$ sinusoidal units, DG size =1024=1024 with top‐$k$ sparsity $k=5\%$, CA3/CA1 MLP widths $(512,256)$ and $(512,256,128)$ respectively, and replay buffer of 200 samples per task. Models are trained with Adam for 10 epochs per CIFAR‐10 task (20 for Tiny‐ImageNet), $\lambda_{\mathrm{ewc}}=0.1$, and contrastive margin $m=0.2$.
Table 1 reports Task‑IL accuracy, Class‑IL accuracy, and inference FLOPs per forward pass for both Small and Large HiCL variants alongside a suite of baselines. We fix the buffer at 200 or 500 samples per task and evaluate on Split CIFAR‑10 (columns 4–6) and Split Tiny‑ImageNet (columns 7–9). Baselines are grouped by methodology: regularization‑based, replay‑based, architectural/hybrid, and the joint offline upper bound.